# LLM-based Agents Suffer from Hallucinations: A Survey of Taxonomy, Methods, and Directions

### v1

@@ Abstract.
LLM-based Agents, Hallucinations, Trustworthiness
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: General and reference Surveys and overviewsccs: Security and privacy AI agent
1. Introduction
@@
Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023a, b; Grattafiori et al., 2024; Guo et al., 2025b; Zhao et al., 2023c) have recently showcased extraordinary capabilities across a broad spectrum of tasks, including language generation (Dong et al., 2019; Wang et al., 2023b), intent comprehension (Geng et al., 2024; Chen et al., 2024c), and knowledge reasoning (Luo et al., 2023; Huang and Chang, 2022). These capabilities are largely attributed to the vast scale of training data (Yu et al., 2023), model architecture innovations (Vaswani et al., 2017), and emergent abilities (Wei et al., 2022a) that arise during instruction tuning (Longpre et al., 2023) and in-context learning (Dong et al., 2022). Building on these breakthroughs, LLM-based Agents (Yao et al., 2023b; Shinn et al., 2023; Zhuang et al., 2023; Chen et al., 2024d; Wu et al., 2024a; Wang et al., 2024f) are becoming increasingly proficient in task automation across a wide range of fields, marking a critical milestone on the path toward Artificial General Intelligence (AGI) (Goertzel, 2014). Moreover, these agents can be orchestrated into a LLM-based Multi-agent System (MAS) (Park et al., 2023; Li et al., 2023b; Hong et al., 2024; Chan et al., 2023; Guo et al., 2024a; Lu et al., 2024), where different agents with distinct specializations collaborate and interact to solve real-world and complex problems beyond the capacity of any single agent through mutual cooperation, such as knowledge sharing (Ba et al., 2024; Han et al., 2024) and collaborative coordination (Zhang et al., 2023d; Zhu et al., 2025a).
@@
Despite the impressive performance achieved by LLM-based agents, their rapid advancement has introduced a new spectrum of safety challenges (Zhang et al., 2024d; Yuan et al., 2024a; Tian et al., 2023). Among these challenges, the problem of Agent Hallucinations represents a particularly significant threat (Liu et al., 2025d; Zhang et al., 2024b; Deng et al., 2025). Previous studies on hallucinations have primarily focused on the field of natural language generation (NLG) (Reiter and Dale, 1997; Gatt and Krahmer, 2018; Dong et al., 2022), where hallucinations refer to the phenomenon in which NLG models generate unfaithful or nonsensical text. Ji et al. (2023a) review recent advances in addressing hallucinations across various NLG tasks. Meanwhile, Huang et al. (2025d) specifically focus on the causes of LLM hallucinations. In this context, hallucinations in LLMs are categorized into factuality and faithfulness hallucinations. Factuality hallucinations highlight discrepancies between the generated content and verifiable real-world facts (Min et al., 2023); while faithfulness hallucinations refer to deviations from the user’s original input (Fabbri et al., 2021). In addition, the authors discuss several effective techniques for detecting and mitigating hallucination issues (Liu et al., 2025g).
@@
Unlike the aforementioned settings, the LLM-based agent is a more sophisticated intelligent system, equipped with goal-directed reasoning and action-taking capabilities. Such agents typically comprise three core modules: brain, perception, and action (Xi et al., 2025; Deng et al., 2025). The brain module is primarily responsible for storing memory and knowledge, supporting reasoning and decision-making for executing tasks; the perception module extends the agent’s perceptual space, enabling it to handle multi-modal environmental inputs; and the action module expands the agent’s action space, allowing it not only to generate textual outputs but also to invoke tools for completing more complex tasks. Therefore, in LLM-based agents, hallucinations are not ”linguistic errors”, but rather a broad category of fabricated or misjudged ”human-like behaviors” that may occur at any stage of the agent’s pipeline. Accordingly, the manifestations and causes of agent hallucinations are considerably more complex. This complexity is reflected in three key aspects: (1) More Diverse Types: Rather than the straightforward response errors of a single model, agent hallucinations are compound behaviors arising from interactions among multiple modules, resulting in a broader and more varied range of hallucination types. (2) Longer Propagation Chains: The aforementioned hallucinations are mostly localized and single-step errors, whereas agent hallucinations often span multiple steps and involve multi-state transitions. Such hallucinations are not limited to the final output; they may also arise during intermediate processes such as perception and reasoning, where they can propagate and accumulate over time. (3) More Severe Consequences: Agent hallucinations involve ”physically consequential” errors, where incorrect embodied actions can directly affect task execution, system devices, and user experiences in the real world. As a result, the cost and risk associated with agent hallucinations are significantly higher. However, existing reviews of LLM-based agents primarily focus on architectural designs and practical applications, while giving far insufficient attention to the importance and urgency of agent hallucinations.
@@
  • First Survey. To the best of our knowledge, this is the first survey to review hallucination issues in LLM-based agents. Moreover, it encompasses recent research on both mitigation and detection methods, offering a broad perspective on the development of LLM-based agents.
@@
  • Open Resource. We provide a well-curated collection of resources, encompassing over 300+ Related Papers, which we make publicly available on GitHub [^2] https://github.com/ASCII-LAB/Awesome-Agent-Hallucinations. to foster community engagement and collaboration.
Figure 1. An overview of agent goal completion. Within the loop, the LLM-based agent carries out external behaviors such as reasoning, execution, perception, and memorization, guided by its internal belief state. Throughout this process, the environment dynamically evolves in response to the agent’s decisions, while task allocation within the LLM-based multi-agent system including broadcasting and structure evolution further enhances the fulfillment of user requirements.
Paper Organization. The remainder of this paper is structured as follows. Section 2 introduces the formal definition of LLM-based agents. Based on this definition, Section 3 presents a new taxonomy of agent hallucinations. Section 4 reviews existing solutions, with particular emphasis on methodologies for hallucination mitigation. Section 5 outlines future research directions. Finally, Section 6 concludes the survey.
2. Formal Definition of LLM-based Agents
2.1. Interaction Dynamics
The interaction dynamics between the agent and the learning environment is usually formulated as a Partially Observable Markov Decision Process (POMDP) (Åström, 1965; Hauskrecht, 2000) which is defined as a 8-tuple: $\mathcal{E}=(\mathcal{S},\mathcal{A},T,\mathcal{G},\mathcal{O},Z,R,\gamma)$. The description of each element in $\mathcal{E}$ is given here:
@@ 2.1. Interaction Dynamics
  • $T:\mathcal{S}\times\mathcal{A}\to P(\mathcal{S})$ is the state transition probability function. For each state-action pair $(s,a)$, it specifies a probability distribution over possible subsequent states. For example, executing a ”turn off lights” action may result in reaching the desired state with 90% probability, while network latency or system errors could cause the state to remain unchanged with 10% probability.
  • $\mathcal{G}$ is the goal space. Each goal $g\in\mathcal{G}$ specifies a particular objective of users (e.g., ”Please help me plan my upcoming trip.”).
@@
This POMDP setting brings together the key challenges agents encounter in open-world settings. For instance, a home assistant agent must interpret multi-modal inputs, navigate uncertainty due to partial observability and randomness, and manage the trade-off between accomplishing immediate tasks and maintaining long-term service excellence.
2.2. LLM-based Agent Loop
@@
    (5) $o_{t+1}=Z(s_{t+1},b_{t},a_{t}).$
    Figure 2. A taxonomy of agent hallucinations with corresponding triggering causes.
@@ 2.3. LLM-based Multi-agent System
Different from the single-agent setting, each LLM-based agent in the MAS must communicate with other agents to accomplish goals. To model these processes, a communication structure $G_{t}$ is introduced among $N$ agents. Furthermore, $G_{t}$ would evolve with time to align with the dynamic adjustments of MAS. Therefore, compared with Section 2.2, the loop of LLM-based MAS includes two additional procedures: Broadcasting and Structure Evolution.
    (8) $p_{t}^{i}=\Phi^{i}(b_{t}^{i},g).$
    (9) $a_{t}^{i}=E^{i}(b_{t}^{i},p_{t}^{i}).$
    (10) $\big\{q_{t}^{i\rightarrow j}\ \big|\ j\in N(i)\big\}=B(b_{t}^{i},p_{t}^{i},G_{t}).$
    (11) $r_{t}=R\left(s_{t},\{a_{t}^{i}\}_{i=1}^{N}\right).$
    (12) ${\rm Pr}\left(s_{t+1}\big|\ s_{t},\{a_{t}^{i}\}_{i=1}^{N}\right)=T\left(s_{t},\{a_{t}^{i}\}_{i=1}^{N}\right).$
    (13) $o_{t+1}^{i}=Z^{i}\big(s_{t+1},b_{t}^{i},a_{t}^{i},\big\{q_{t}^{j\rightarrow i}\ \big|\ j\in N(i)\big\}\big),$
    (14) $m_{t+1}^{i}=L_{M}^{i}\left(m_{t}^{i},a_{t}^{i},\big\{q_{t}^{j\rightarrow i}\ \big|\ j\in N(i)\big\},o_{t+1}^{i}\right).$
  • Belief Update: Then the agent $i$ refines its belief state as follows,
    (15) $b_{t+1}^{i}=L_{B}^{i}\left(b_{t}^{i},m_{t+1}^{i},a_{t}^{i},\big\{q_{t}^{j\rightarrow i}\ \big|\ j\in N(i)\big\},r_{t}^{i},o_{t+1}^{i},g\right).$
    (16) $G_{t+1}=U\left(G_{t},\{b_{t+1}^{i}\}_{i=1}^{N}\right).$
Fig. 1 provides an overview of agent goal completion, which incorporates the agent loop, environment transitions, and the collaborative process in LLM-based MAS.
@@
Definition of Agent hallucinations According to the proposed internal–external distinction, agent hallucinations refer to fictitious or erroneous operations in LLM-based agents, driven by overconfident “human-like behaviors” in reasoning, execution, perception, memorization, and communication. Unlike LLM hallucinations, which are typically limited to linguistic errors, agent hallucinations go beyond language generation, manifesting as compound deviations arising from the coupling and interaction of multiple modules. These deviations can escalate to task failures and even physical safety risks. In contrast to other types of agent errors, agent hallucinations are often spontaneously generated within the agent’s own cognitive processes, representing an endogenous error form. Furthermore, owing to their cross-module propagation and hidden cognitive nature, agent hallucinations pose an especially serious threat, making them a central challenge in the pursuit of building safe and reliable agents.
3. Taxonomy of Agent Hallucinations
@@
We first provide a formal definition of agent hallucinations. Based on this definition, we then introduce the new taxonomy that includes the following five types of agent hallucinations: Reasoning Hallucinations, Execution Hallucinations, Perception hallucinations, Memorization Hallucinations, and Communication Hallucinations, as illustrated in Fig. 2. Detailed descriptions of each type of agent hallucinations are provided in the following sub-sections, and representative hallucination examples are presented in Fig. 3.
3.1. Reasoning Hallucinations
@@
Reasoning serves as the cornerstone of an agent’s functionality, influencing behavior analysis and decision-making (Cao et al., 2025; He et al., 2025a; Zhang et al., 2025f). Upon receiving a specific goal $g$, the agent first leverages its own reasoning capabilities to perform goal understanding for inferring the user’s true intention:
@@
    (17) $\mathcal{I}=\text{Understand}(b_{0},g),$
@@
  • Intention Decomposition: Typically, there are two decomposition methods: pre-defined decomposition and dynamic decomposition (Ephrati and Rosenschein, 1994). In the first one, these sub-intentions are specified in advance:
    (18) $\{\mathcal{I}_{t}\}_{t=0}^{n}=\text{Pre-decompose}(b_{0},\mathcal{I}).$
@@
    (19) $\mathcal{I}_{t}=\text{Dyn-decompose}(b_{t},\mathcal{I}).$
    The decomposition results are dynamically optimized based on the current belief state $b_{t}$, enabling timely and continuous refinement for the subsequent workflow (Sapkota et al., 2025; Liu et al., 2025e; Yuan et al., 2025).
@@
    (20) $p_{t}=\text{Plan}(b_{t},\mathcal{I}_{t}).$
Definition of Reasoning Hallucinations Reasoning hallucinations refer to the phenomenons where LLM-based agents generate seemingly plausible plans which are in fact logically flawed or unsupported. Reasoning hallucinations occur in either of above three stages, and we categorize them into three types: Goal Understanding Hallucinations, manifested as misinterpretation of goal semantics, typically caused by Problematic Objective Expression and Inadequate Subjective Comprehension; Intention Decomposition Hallucinations, characterized by the flawed decomposition processes that produce irrelevant or infeasible sub-intentions, often due to Deficient Dependency Modeling; Planning Generation Hallucinations represent hallucinations that occur when an agent generates each specific plan of a sub-intention, typically caused by the Misinterpretation and Misapplication of Planning Information.
@@ 3.1.1. Problematic Objective Expression
Understanding the user’s true intention from $g$ is the first and essential step for effective reasoning (Zhang et al., 2025a). Objectively, when the expression of goal information carries a certain degree of semantic vagueness, it can easily lead to erroneous parsing of user intention and induce reasoning hallucinations (Zheng et al., 2023; Linok et al., 2024; Zhang et al., 2025b). This semantic vagueness can be mainly attributed to two issues: Incomplete Goal Specification and Ambiguous Content Presence (Liu et al., 2023c; Keluskar et al., 2024). To address semantic vagueness in user input, a plausible and effective approach is to endow the agent with the capability of active clarification (Deng et al., 2023; Kobalczyk et al., 2025; Chen et al., 2025c). By engaging in appropriate proactive interactions with the user, the agent can gain a deeper understanding of the user’s true needs and intentions.
3.1.2. Inadequate Subjective Comprehension
@@
In addition to above objective issues, there also exist subjective limitations in the agent’s comprehension capability. These are mainly manifested as Instruction-following Deviation (Verma et al., 2025; Han et al., 2025a), as well as Long-range Contextual Misuse (Xu and Wang, 2025). Specifically, instruction following refers to an agent’s ability to adhere to user-provided instructions in order to accurately comprehend intentions (Abbasian et al., 2023). When this procedure breaks down, the agent struggles to segment critical fields and extract key information from instructions, leading to distorted understanding and reasoning hallucinations (Huang et al., 2024a; Wu et al., 2025a; Shi et al., 2024b). Additionally, when agents handle excessively long input goals, they may fail to capture long-range contexts and instead over-rely on the most recent tokens, resulting in contextual misuse (Zhao et al., 2023a; Chu et al., 2024a; Zhu et al., 2024; Yu and McQuade, 2025). Such deficiencies cause agents to violate explicitly stated user preferences (Lai et al., 2024b; Kondapaneni et al., 2024).
Figure 3. The depiction of different types of agent hallucinations, each illustrated with a representative example.
3.1.3. Deficient Dependency Modeling
@@
Intention decomposition is a crucial phase where agents determine the successive reasoning process via the decomposed sub-intentions (Wang et al., 2025g; Tudor et al., 2025; Zhao et al., 2025a). These sub-intentions typically form a sequentially dependent chain, where the completion of each sub-intention pre-supposes the successful fulfillment of its predecessors. Inadequate modeling of dependency relationships among these sub-intentions can give rise to three types of errors: Sub-intention Omission, Sub-intention Redundancy (Wan et al., 2024b; Li et al., 2024a) and Sub-intentions Disorder. All three severely compromise reasoning integrity and efficiency, ultimately leading to reasoning hallucinations (Wang et al., 2025h). Sub-intention omission occurs when agents fail to identify essential sub-intentions, leading to the absence of critical reasoning steps (Qiu et al., 2025). Sub-intention redundancy arises when agents erroneously introduce task-irrelevant sub-intentions, causing reasoning processes to deviate from the core intention and generating hallucinatory content (Berijanian et al., 2025; Yu et al., 2025; Qiu et al., 2025). Sub-intention disorder refers to the phenomenon in which an agent arranges sub-intentions with a sequential relationship in the wrong order during reasoning, causing it to fail to obtain the results of prerequisite sub-intentions, thereby missing key information and further generating hallucinations (Wang et al., 2023a).
3.1.4. Planning Information Misinterpretation
@@
In Eq.(20), the agent relies on two key sources of information to generate sub-intention plans: Operable Objects and Self-knowledge. For operable objects, the agent must recognize both explicit and implicit relational connections between entities (e.g., spatial, temporal, or social relationships) (Chu et al., 2025; Huang et al., 2024b; Zuo et al., 2025). The misinterpreting information related to these operable objects can lead the agent to construct plans based on incorrect assumptions, thereby inducing planning generation hallucinations (Manakul et al., 2023; Xu et al., 2025b). With respect to self-knowledge, the agent is expected to plan reasonably within the bound of its own self-knowledge to avoid planning errors. When confronted with planning problems beyond its knowledge boundary, the agent tends to respond with excessive confidence, generating answers that sound certain but are actually incorrect (Liu et al., 2025b; Xiong et al., 2023; Deng et al., 2024c).
3.1.5. Planning Information Misapplication
@@
Even when an agent correctly identifies the operable objects within its knowledge boundary, its ability to effectively utilize them remains constrained by its Logical Reasoning and Introspection Capacity (Mirzadeh et al., 2024). Logical fallacies primarily manifest as rule conflicts and causal misattributions (Liu et al., 2025a). Similar to humans, when agents encounter issues such as rule conflicts, causal inversion, or misattribution during causal reasoning (Wu and Cardie, 2025; Berglund et al., 2023), they often fail to complete the reasoning process or resolve underlying logical paradoxes, ultimately leading to erroneous planning (Choi et al., 2024; Xiao et al., 2024; Xu et al., 2025d). Additionally, introspection is one of the key mechanisms by which agents achieve learning, cognition, and knowledge updating (Shinn et al., 2023), agents that lack sufficient introspection capacity are unable to effectively improve their performance by revisiting logical fallacies in their reasoning (Belle et al., 2025; Renze and Guven, 2024). Such agents may misinterpret feedback signals and accurately identify the sources of error (Liang et al., 2023), thereby making misguided adjustments to their reasoning processes (Zhang et al., 2024g). Even when feedback is correctly interpreted, the introspection process may still be flawed: agents might wrongly judge their reasoning as consistent despite existing logical contradictions, or mistakenly believe they have considered all relevant factors while overlooking critical variables (Yan et al., 2024; Li et al., 2024e).
@@
Definition of Execution Hallucinations Execution hallucinations refer to the phenomenon where LLM-based agents claim to have completed certain sub-stages during the execution phase, but in reality, they have not actually been performed or accomplished. According to the above two sub-stages, we divide execution hallucinations into two types: Tool Selection Hallucinations, where non-existent or irrelevant tools are chosen overconfidently; and Tool Calling Hallucinations, involving incorrect, omitted, or fabricated parameter fillings.
3.2. Execution Hallucinations

### v2

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
@@
LLanguage Models (LLMs) [achiam2023gpt, touvron2023llama, touvron2023llama2, grattafiori2024llama, guo2025deepseek] have recently showcased extraordinary capabilities across a broad spectrum of tasks, including language generation [wang2023grammar], intent comprehension [geng2024large], and knowledge reasoning [luo2023reasoning]. These capabilities are largely attributed to the vast scale of training data [yu2023large], model architecture innovations [vaswani2017attention], and emergent abilities [wei2022emergent] that arise during instruction tuning [longpre2023flan] and in-context learning [dong2022survey]. Building on these breakthroughs, LLM-based Agents [yao2023react, shinn2023reflexion, chen2024agent, wu2024graph, wang2024survey] are becoming increasingly proficient in task automation across a wide range of fields, marking a critical milestone on the path toward Artificial General Intelligence (AGI) [goertzel2014artificial]. Moreover, these agents can be orchestrated into a LLM-based Multi-agent System (MAS) [park2023generative, li2023camel, hong2024metagpt, chan2023chateval, guo2024large], where different agents with distinct specializations collaborate and interact to solve real-world and complex problems beyond the capacity of any single agent through mutual cooperation, such as knowledge sharing [ba2024cautiously] and collaborative coordination [zhu2025multiagentbench].
@@
Despite the impressive performance achieved by LLM-based agents, their rapid advancement introduces a spectrum of safety challenges [zhang2024agent, yuan2024r, tian2023evil]. Among these challenges, Agent Hallucinations represents a particularly significant threat [liu2025advances, r55, deng2025ai]. Previous studies on hallucinations have primarily focused on the field of natural language generation (NLG) [dong2022survey], where hallucinations refer to the phenomenon in which NLG models generate unfaithful or nonsensical text. Ji et al. review recent advances in addressing hallucinations across various NLG tasks [ji2023survey]. Meanwhile, Huang et al. specifically focus on the causes of LLM hallucinations [huang2025survey]. In this context, hallucinations in LLMs are categorized into factuality and faithfulness hallucinations. Factuality hallucinations highlight discrepancies between the generated content and verifiable real-world facts [min2023factscore]; while faithfulness hallucinations refer to deviations from the user’s original input [fabbri2021qafacteval]. In addition, the authors discuss several effective methods for detecting and mitigating hallucination issues [liu2025reducing].
@@
Unlike the aforementioned settings, the LLM-based agent is a more sophisticated intelligent system, equipped with goal-directed reasoning and action-taking capabilities. Such agents typically comprise three core modules: brain, perception, and action [xi2025rise]. The brain module is primarily responsible for storing memory and knowledge, supporting reasoning and decision-making for executing tasks; the perception module extends the agent’s perceptual space, enabling it to handle multi-modal environmental inputs; and the action module expands the agent’s action space, allowing it not only to generate textual outputs but also to invoke tools for completing more complex tasks. Therefore, in LLM-based agents, hallucinations are not ”linguistic errors”, but rather a broad category of fabricated or misjudged ”human-like behaviors” that may occur at any stage of the agent’s pipeline. Accordingly, the manifestations and causes of agent hallucinations are considerably more complex. This complexity is reflected in three key aspects: 1) More Diverse Types: Rather than the straightforward response errors of a single model, agent hallucinations are compound behaviors arising from interactions among multiple modules, resulting in a broader and more varied range of hallucination types. 2) Longer Propagation Chains: The aforementioned hallucinations are mostly localized and single-step errors, whereas agent hallucinations often span multiple steps and involve multi-state transitions. Such hallucinations are not limited to the final output; they may also arise during intermediate processes such as perception and reasoning, where they can propagate and accumulate over time. 3) More Severe Consequences: Agent hallucinations involve ”physically consequential” errors, where incorrect embodied actions can directly affect task execution, system devices, and user experiences in the real world. As a result, the cost and risk associated with agent hallucinations are significantly higher. However, existing reviews of LLM-based agents primarily focus on architectural designs and practical applications, while giving far insufficient attention to the importance and urgency of agent hallucinations.
Figure 1: An overview of agent goal completion. Within the loop, the LLM-based agent carries out external behaviors such as reasoning, execution, perception, and memorization, guided by its internal belief state. Throughout this process, the environment dynamically evolves in response to the agent’s decisions, while task allocation within the LLM-based multi-agent system including broadcasting and structure evolution further enhances the fulfillment of user requirements.
@@
  • First Survey. To the best of our knowledge, this is the first survey to review hallucination issues in LLM-based agents. It encompasses recent research on both mitigation and detection methods, offering a broad perspective on the development of LLM-based agents.
@@
  • Open Resource. We provide a well-curated collection of resources, encompassing over 200+ related papers, which we make publicly available on GitHub [^2] https://github.com/ASCII-LAB/Awesome-Agent-Hallucinations. to foster community engagement and collaboration.
Paper Organization. The remainder of this paper is structured as follows. Section II introduces the formal definition of LLM-based agents. Based on this definition, Section III presents a new taxonomy of agent hallucinations. Section IV reviews existing solutions, with particular emphasis on methodologies for hallucination mitigation. Section V outlines future research directions. Section VI concludes the survey.
II Formal Definition of LLM-based Agents
Fig. 1 provides an overview of agent goal completion. The formal definition of LLM-based agents including interaction dynamics and LLM-based agent loop is given here.
II-A Interaction Dynamics
The interaction dynamics between the agent and the learning environment is usually formulated as a Partially Observable Markov Decision Process (POMDP) [aastrom1965optimal] which is defined as a 8-tuple: $\mathcal{E}=(\mathcal{S},\mathcal{A},T,\mathcal{G},\mathcal{O},Z,R,\gamma)$. The description of each element in $\mathcal{E}$ is given here:
@@ II-A Interaction Dynamics
  • $T:\mathcal{S}\times\mathcal{A}\to P(\mathcal{S})$ is the state transition probability function. For each state-action pair $(s,a)$, it specifies a probability distribution over possible subsequent states. For example, executing a “turn off lights” action may result in reaching the desired state with 90% probability, while network latency or system errors could cause the state to remain unchanged with 10% probability.
  • $\mathcal{G}$ is the goal space. Each goal $g\in\mathcal{G}$ specifies a particular objective of users.
@@
II-B LLM-based Agent Loop
@@
    $o_{t+1}=Z(s_{t+1},b_{t},a_{t}).$ (5)
@@
Different from the single-agent setting, in the LLM-based MAS, each agent must communicate with other agents to accomplish goals. To model these processes, a communication structure $G_{t}$ is introduced among $N$ agents. Furthermore, $G_{t}$ would evolve with time to align with the dynamic adjustments of MAS. Therefore, compared with the above loop, the loop of LLM-based MAS is given as “Reasoning–Execution–Broadcasting–Feedback–Environment Transition–Perception–Memorization–Belief Update–Structure Evolution”. This loop includes two additional procedures: Broadcasting and Structure Evolution. Broadcasting means that the agent broadcasts its message to neighboring nodes according to its plan, while structure evolution indicates that the communication structure $G_{t}$ can be updated in each iteration. The complete description of the loop of MAS is given in Appendix A.
@@
Definition of Agent hallucinations According to the proposed internal–external distinction, agent hallucinations refer to fictitious or erroneous operations in LLM-based agents, driven by overconfident “human-like behaviors” in reasoning, execution, perception, memorization, and communication. Unlike LLM hallucinations, which are typically limited to linguistic errors, agent hallucinations go beyond language generation, manifesting as compound deviations arising from the coupling and interaction of multiple modules. These deviations can escalate to task failures and even physical safety risks. In contrast to other types of agent errors, agent hallucinations are often spontaneously generated within the agent’s own cognitive processes, representing an endogenous error form. Furthermore, owing to their cross-module propagation and hidden cognitive nature, agent hallucinations pose an especially serious threat, making them a central challenge in the pursuit of agent safety.
Figure 2: A taxonomy of agent hallucinations. It includes five hallucination types and nine hallucination sub-types with corresponding triggering causes.
III Taxonomy of Agent Hallucinations
@@
We first provide a formal definition of agent hallucinations. Based on this definition, we then introduce the new taxonomy that includes the following five types of agent hallucinations: Reasoning Hallucinations, Execution Hallucinations, Perception hallucinations, Memorization Hallucinations, and Communication Hallucinations, as illustrated in Fig. 2. Detailed descriptions of each type of agent hallucinations are provided in the following sub-sections. Representative hallucination examples are presented in Appendix B.
III-A Reasoning Hallucinations
@@
Reasoning serves as the cornerstone of an agent’s functionality, influencing behavior analysis and decision-making [cao2025large, he2025plan, zhang2025igniting]. Upon receiving a specific goal $g$, the agent first leverages its own reasoning capabilities to perform goal understanding for inferring the user’s true intention:
@@
    $\mathcal{I}=\text{Understand}(b_{0},g),$ (8)
@@
  • Intention Decomposition: Typically, there are two decomposition methods: pre-defined decomposition and dynamic decomposition. In the first one, these sub-intentions are specified in advance:
    $\{\mathcal{I}_{t}\}_{t=0}^{n}=\text{Pre-decompose}(b_{0},\mathcal{I}).$ (9)
@@
    $\mathcal{I}_{t}=\text{Dyn-decompose}(b_{t},\mathcal{I}).$ (10)
    The decomposition results are dynamically optimized based on the current belief state $b_{t}$, enabling timely and continuous refinement for the subsequent workflow [sapkota2025ai].
@@
    $p_{t}=\text{Plan}(b_{t},\mathcal{I}_{t}).$ (11)
Definition of Reasoning Hallucinations Reasoning hallucinations refer to the phenomenons where LLM-based agents generate seemingly plausible plans which are in fact logically flawed or unsupported. Reasoning hallucinations occur in either of above three stages, and we categorize them into three types: Goal Understanding Hallucinations (GUHs), manifested as misinterpretation of goal semantics, typically caused by problematic objective expression and inadequate subjective comprehension; Intention Decomposition Hallucinations (IDHs), characterized by the flawed decomposition processes that produce irrelevant or infeasible sub-intentions, often due to deficient dependency modeling; Planning Generation Hallucinations (PGHs) represent hallucinations that occur when an agent generates each specific plan of a sub-intention, typically caused by the misinterpretation and misapplication of planning information.
@@
1) Problematic Objective Expression. Understanding the user’s true intention from $g$ is the first and essential step for effective reasoning [zhang2025metamind]. Objectively, when the expression of goal information carries a certain degree of semantic vagueness, it can easily lead to erroneous parsing of user intention and induce reasoning hallucinations [zheng2023does, linok2024beyond]. This semantic vagueness can be mainly attributed to two issues: Incomplete Goal Specification and Ambiguous Content Presence [liu-etal-2023-afraid, keluskar2024llms]. To address semantic vagueness in user input, a plausible and effective approach is to endow the agent with the capability of active clarification [kobalczyk2025active, chen2025learning]. By engaging in appropriate proactive interactions with the user, the agent can gain a deeper understanding of the user’s true needs and intentions.
@@
2) Inadequate Subjective Comprehension. In addition to above objective issues, there also exist subjective limitations in the agent’s comprehension capability. These are mainly manifested as Instruction-following Deviation [verma2025graft], as well as Long-range Contextual Misuse [xu2025scrnet]. Specifically, instruction following refers to an agent’s ability to adhere to user-provided instructions in order to accurately comprehend intentions [abbasian2023conversational]. When this procedure breaks down, the agent struggles to segment critical fields and extract key information from instructions, leading to distorted understanding and reasoning hallucinations [huang2024opera, wu2025clarifycoder, shi2024navigating]. Additionally, when agents handle excessively long input goals, they may fail to capture long-range contexts and instead over-rely on the most recent tokens, resulting in contextual misuse [zhao2023large, chu2024beamaggr, zhu2024knowagent]. Such deficiencies cause agents to violate explicitly stated user preferences [lai2024vision].
@@
3) Deficient Dependency Modeling. Intention decomposition is a crucial phase where agents determine the successive reasoning process via the decomposed sub-intentions [wang2025medagent, tudor2025vecsr, zhao2025mactgmultiagentcollaborativethought]. These sub-intentions typically form a sequentially dependent chain, where the completion of each sub-intention pre-supposes the successful fulfillment of its predecessors. Inadequate modeling of dependency relationships among these sub-intentions can give rise to three types of errors: Sub-intention Omission, Sub-intention Redundancy [wan2024cot, li2024focus] and Sub-intentions Disorder. All three severely compromise reasoning integrity and efficiency, ultimately leading to reasoning hallucinations [wang2025tdag]. Sub-intention omission occurs when agents fail to identify essential sub-intentions, leading to the absence of critical reasoning steps [qiu2025co]. Sub-intention redundancy arises when agents erroneously introduce task-irrelevant sub-intentions, causing reasoning processes to deviate from the core intention and generating hallucinatory content [berijanian2025comparative, yu2025dyna, qiu2025co]. Sub-intention disorder refers to the phenomenon in which an agent arranges sub-intentions with a sequential relationship in the wrong order during reasoning, causing it to fail to obtain the results of prerequisite sub-intentions, thereby missing key information and further generating hallucinations [wang2023steps].
@@
4) Planning Information Misinterpretation. In Eq.(11), the agent relies on two key sources of information to generate sub-intention plans: Operable Objects and Self-knowledge. For operable objects, the agent must recognize both explicit and implicit relational connections between entities (e.g., spatial, temporal, or social relationships) [chu2025llm+, huang2024targa, zuo2025kg4diagnosis]. The misinterpreting information related to these operable objects can lead the agent to construct plans based on incorrect assumptions, thereby inducing planning generation hallucinations [manakul2023selfcheckgpt]. With respect to self-knowledge, the agent is expected to plan reasonably within the bound of its own self-knowledge to avoid planning errors. When confronted with planning problems beyond its knowledge boundary, the agent tends to respond with excessive confidence, generating answers that sound certain but are actually incorrect [liu2025uncertainty].
@@
5) Planning Information Misapplication. Even when an agent correctly identifies the operable objects within its knowledge boundary, its ability to effectively utilize them remains constrained by its Logical Reasoning and Introspection Capacity [mirzadeh2024gsm]. Logical fallacies primarily manifest as rule conflicts and causal misattributions [liu2025enhancing]. Similar to humans, when agents encounter issues such as rule conflicts, causal inversion, or misattribution during causal reasoning [wu2025reasoning], they often fail to complete the reasoning process or resolve underlying logical paradoxes, ultimately leading to erroneous planning [xu2025mragent]. Additionally, introspection is one of the key mechanisms by which agents achieve learning, cognition, and knowledge updating [shinn2023reflexion], agents that lack sufficient introspection capacity are unable to effectively improve their performance by revisiting logical fallacies in their reasoning [belle2025agents]. Such agents may misinterpret feedback signals and accurately identify the sources of error [liang2023encouraging], thereby making misguided adjustments to their reasoning processes [zhang2024self]. Even when feedback is correctly interpreted, the introspection process may still be flawed: agents might wrongly judge their reasoning as consistent despite existing logical contradictions, or mistakenly believe they have considered all relevant factors while overlooking critical variables [li2024hindsight].
@@
Definition of Execution Hallucinations Execution hallucinations refer to the phenomenon where LLM-based agents claim to have completed certain sub-stages during the execution phase, but in reality, they have not actually been performed or accomplished. According to the above two sub-stages, we divide execution hallucinations into two types: Tool Selection Hallucinations (TSHs), where non-existent or irrelevant tools are chosen overconfidently; and Tool Calling Hallucinations (TCHs), involving incorrect, omitted, or fabricated parameter fillings.
III-B Execution Hallucinations